12 ml of a 80% alcohol solution was mixed with 8 ml of a 40% alcohol solution. Find the concentration of the new mixture.
Answer by math_tutor2020(3817)


Bottle A = 12 ml of 80% alcohol
Bottle B = 8 ml of 40% alcohol

Bottle A has 12*0.80 = 9.6 ml of pure alcohol
Bottle B has 8*0.40 = 3.2 ml of pure alcohol
Combined, we have 9.6+3.2 = 12.8 ml of pure alcohol out of 12+8 = 20 ml total solution.

The alcohol concentration of the mix of A and B is 12.8/20 = 0.64 = 64%

Answer: 64% alcohol solution